Please help me to solve!!!
"Woof, Woof, Meow!"
Waggles is tied to a chain and has a playground of 154 square feet. The neighbor's cat Fluffy is currently 2 1/2 yards from the tie-down post. Should Fluffy be concerned?
Thanks for helping :)

Answer by stanbon(75887)

Waggles is tied to a chain and has a playground of 154 square feet. The neighbor's cat Fluffy is currently 2 1/2 yards from the tie-down post. Should Fluffy be concerned?
------
Find radius of Waggles' playground:
pi*r^2 = 154 sq. ft
r^2 = 154/pi
r = 7 ft
-----
Distance of Fluffy from the center of Waggles' playground:
2 1/2 yds = (5/2)(3 ft) = 15/2 = 7.5 feet.
------
Fluffy does not have to be concerned since he/she is outside
the radius of Waggles" playground.
--------------------
Cheers,
Stan H.
====================